# Johannes Matzat        Aiko Schmeißer

### v4

@@ 5.2 Addressing Identification Challenges
DiD-RDD. We continue presenting results for our RDD-motivated tests to probe the validity of the underlying parallel trends assumption of the DiD model. [^30] One particular concern for the parallel trends assumption would arise if union elections were endogenously timed around federal election dates. Appendix Figure A.4 investigates whether union elections follow political cycles. Across years with and without federal elections, there are no strong differences in the number of union elections held and the probability of winning a union election, in particular not around the week of federal elections. Thus, we do not see evidence that employers or unions successfully manipulate union election dates to change union support around federal election cycles. Figure 3 focuses on the partisan composition of contributions, while effects on the total amount of contributions are presented in Appendix Figure A.5. Results are always reported separately for workers and managers. We first analyze the heterogeneous effects of unionization across the vote share distribution. Panel (a) of Figure 3 displays the $\delta_{g}$ coefficients from model (3) for the interaction between the post-election dummy and different vote share categories. The results show that there are no significantly different trends among losing elections with a vote share of 0-20% or 35-50% relative to those with 20-35%, for contributions from both workers and managers. The partisan composition of contributions thus appears to evolve similarly across losing establishments with different vote shares, suggesting that unobserved shocks correlated with voting results are unlikely to drive our results. [^31] In Appendix Figure A.6, we also investigate whether pre-trends in the contribution composition are similar across the vote share distribution. For this, we estimate the following modified version of model (3): $y_{jk}=\alpha_{j}+\beta_{kg_{j}}+\sum_{g}\delta^{PRE}_{g}\times\Big{(}\mathbbm{1}[k<-1]\times\mathbbm{1}[V_{j}\in\nu^{g}]\Big{)}+\sum_{g}\delta^{POST}_{g}\times\Big{(}\mathbbm{1}[k\geq 0]\times\mathbbm{1}[V_{j}\in\nu^{g}]\Big{)}+\epsilon_{jk}$ (6) The results show that none of the $\delta^{PRE}_{g}$ coefficients are significantly different from zero, indicating that, also before the union election, contribution patterns evolved similarly across establishments with different voting results. Moreover, the results indicate whether treatment effects are heterogeneous across vote shares among winning union elections. For the composition of contributions from managers, the estimate is significant across all vote share categories above 50%. Thus, the political response of managers does not appear to depend on whether workers won the union election with large or small margins of victory. For workers, the effect on partisan support is significant only for vote shares between 50 and 80% and appears smaller for elections won by a large margin.
@@
Panel (b) of Figure 3 presents coefficients from the DiD model (1) when restricting the sample to establishments with increasingly close election results. Establishments with more similar voting results can be expected to be more similar in other characteristics and to be exposed to more similar shocks, which makes the parallel trends assumption more plausible. Results are reported in 5% steps of the union vote share bandwidth around the 50% cutoff. Our baseline results from Figure 2 include only elections with a pro-union vote share between 20 and 80%, i.e., a bandwidth of 30%. Figure 3 shows that treatment effects are very similar when instead using all elections. More importantly, the results are also very stable when focusing on closer elections. Even when restricting the sample to establishments that won with a maximum vote margin of 5%, we see a positive and significant effect on the composition of campaign contributions for managers. Similarly, for workers a maximum vote margin of 10% already yields a positive and significant effect.
DiD-IV. We also assess the sensitivity of our DiD results when exploiting arguably exogenous variation in unionization from shocks to the salience of workplace safety before the union election. Table 3 reports the results of our DiD-IV approach. The first-stage results show that sector-level fatal work accidents are a significant predictor of the union election outcome, with an F-statistic of 16.5. We find that the positive effect of spikes in work accidents on unionization is stronger in sectors where work accidents are more common, i.e., where workplace safety may be a greater concern for workers. The second-stage results confirm our main findings from the DiD model, highlighting a leftward shift in campaign contributions in response to unionization. [^32] We also report results when estimating model (5) by OLS. Given that the outcome is the change in outcomes before versus after the election, the results are very similar to those obtained from our main DiD model (1). Small differences arise from the inclusion of additional controls in model (5) and the exclusion of the cycle in which the union election takes place. The magnitude of the coefficients is comparable but slightly larger than in the DiD model. As compliers respond to information on fatal work accidents, we deem it plausible that they also react more strongly to information provided by unions and to changes to their work environment induced by unionization. The estimates are considerably less precise, however. While the effects on the party composition of contributions from managers are still significant at the 5% level, the effects for workers are no longer significant. We thus use our DiD-IV approach to validate the main results and proceed with the DiD model for the analysis of mechanisms in Section 6.333333We also verify the DiD-IV approach with a falsification exercise. We re-estimate model (5) using the change in campaign contribution patterns between $t-1$ and $t-2$ as the outcome. We do not find any evidence for pre-existing differential trends related to spikes in fatal work accidents.
@@ 5.3 Robustness
We now discuss further robustness checks for our main DiD estimates. Results are presented in Appendix Tables A.4 to A.9.
Extensive versus intensive margin weights. The percentage effect interpretation of the IHS results implicitly assumes a weight of extensive and intensive margin changes in contribution amounts (Chen and Roth, 2024). To understand the role of both margins, we perform two analyses. First, in Appendix Table A.4, we directly study extensive margin effects using as outcome variable a dummy for any positive donation by an employee at the establishment. The DiD estimates show no significant effect of unionization on the likelihood that at least one employee in the establishment donates a positive amount to any candidate. However, for the partisan composition, we find that unionization results in an increased likelihood of employees donating to Democratic relative to Republican candidates of 2.8 percentage points. This effect is observed for both workers’ and managers’ donations and suggests that extensive margin changes may explain at least some of the partisan shift in campaign contributions.
Second, we analyze the robustness of our main results when placing explicit weights on extensive and intensive margin changes. For that, we follow the suggestion in Chen and Roth (2024) and use the following outcome transformation:
$m(\$)=\begin{cases}\ln(\$)&\text{for }\$>0\\ -x&\text{for }\$=0.\end{cases}$ (7)
This transformation gives equal weight to the extensive margin effect of moving from 0 to 1 and to the intensive margin effect of increasing donations by 100 log points for establishments with positive donations. Appendix Table A.5 reports results for alternative weights $x\in\{0,0.1,0.7,1,3\}$. Setting $x=0.7$, we find the same DiD estimates as for the IHS-transformed outcomes, which indicates that our main results implicitly weight an extensive margin change from 0 to 1 as equivalent to an intensive margin change in donations by 70 log points. Importantly, the results remain similar, growing only slightly in magnitude, for alternative weights of 10, 100, or 300 log points. For $x=0$, which shuts off extensive margin changes by treating zero amounts as equal to one USD, we also find very similar results as for the weighting implied by the IHS specification.
@@
Outcome transformations.Roth and Sant’Anna (2023) point out that the parallel trends assumption of a DiD design generally implies a functional form restriction on potential outcomes. Different transformations of the outcome may imply different parallel trends assumptions. We therefore test the sensitivity of our results to alternative outcome transformations. In addition to transformations using IHS and equation (7), we consider the log function after adding one to the amounts, take the quartic root of amounts [^34] As shown by Thakral and Tô (2023), power transformations – unlike quasi-logarithmic transformations (such as the IHS or log(y+1) function) – maintain scale invariance of the implied semi-elasticities., as well as leave amounts untransformed (in 2010 USD). Results, shown in Panels B, C, and D of Appendix Table A.6, yield the same conclusions as the results for the IHS-transformed outcomes.
@@
Staggered DiD estimators. The recent econometrics literature has proposed different methods to circumvent issues of treatment effect heterogeneity in staggered DiD designs. All the proposed estimation strategies have in common that they restrict the set of effective comparison units by ruling out the use of early-treated units in the estimation of treatment effects for currently-treated units. They differ, however, in terms of how exactly comparison units are identified and used in the estimation, as well as in terms of how cohort- or individual-specific treatment effect estimates are aggregated. [^35] In our stacking approach of model (1), we effectively only compare winning elections to losing elections that were held in the same period, i.e., we only use never-treated units in the comparison group. The strategies by Borusyak et al. (forthcoming) and Callaway and Sant’Anna (2021), in contrast, also allow including not-yet-treated units in the comparison group. Both approaches differ in that Borusyak et al. (forthcoming) use the average pre-treatment outcome over all pre-treatment periods, whereas Callaway and Sant’Anna (2021) only use the outcome one period before treatment start. In terms of aggregation, Gardner (2021) shows that the stacking approach identifies a convexly weighted average of cohort-specific treatment effects where the weights are given by the number of treated units and the variance of treatment within each cohort. In comparison, Borusyak et al. (forthcoming) and Callaway and Sant’Anna (2021) first estimate unit- or cohort-specific effects and then aggregate through a simple average across treated units. Callaway and Sant’Anna (2021) also allow other weights, but we use the default option where cohort-specific estimates are weighted by the number of treated units in each cohort. In Panels E and F of Appendix Table A.6, we present results from the imputation approach of Borusyak et al. (forthcoming) and the estimator developed by Callaway and Sant’Anna (2021). The estimates are very similar to our stacked DiD results.
@@
Sample restrictions. For our baseline results, we have restricted the sample to establishments with at least one matched contribution over the observation period. If employees have not contributed before the union election and start contributing in response to unionization, concerns of sample selection bias may arise. In Appendix Table A.2, we show that there are no effects of unionization on the likelihood of observing any contribution after the union election (in line with estimated null effects on total contribution amounts). Nevertheless, Panel G of Appendix Table A.6 also shows results when restricting the analysis to establishments that have at least one matched contribution in any cycle before the union election. This reduces our sample size but leaves results largely unchanged. Another concern may arise from selected establishment survival (Frandsen, 2021; Wang and Young, 2023). In particular, we may wrongly assign zero contribution amounts to establishments that do not yet exist or that have already been closed. In order to restrict the analysis to a window of known survival, we exclude for each establishment the cycles before the first matched contribution and the cycles after the last matched contribution. As shown in Panel H, this restriction increases the size of our estimates suggesting that our baseline results can be interpreted as lower bounds.
@@
Manager-worker classifications. In Appendix Table A.7, we check whether our results are sensitive to the exact definition of managers and supervisors versus non-managerial workers. To see whether the political response is different for lower- and upper-tier managers, we use more stringent definitions of managers/supervisors. First, we vary the cutoff for the importance of supervisor tasks (Panels B and C). Second, we only consider “Management Occupations” (SOC group 11) and treat all other occupations (including those with high importance of supervisor tasks) as workers (Panel D). The results do not change much with these alternative classifications. Even for more upper-tier managers unionization leads to an increase in the support for Democrats relative to Republicans.
@@
Federal versus local candidates. Next, we examine whether our effects are limited to contributions to candidates for either federal or local (i.e., state) offices. Focusing on federal races offers the advantage of analyzing a set of comparable candidates who are potentially relevant to citizens nationwide. At the same time, U.S. legislation on labor issues, which unions may focus on when endorsing candidates, is enacted not only at the federal level but also at the state level (e.g., state-specific minimum wages, right-to-work laws). Appendix Table A.8 shows that our estimates are driven by contributions to both federal and local candidates. The effect sizes are somewhat larger for contributions to candidates running for federal office, but at both levels, we observe a significant shift in donations from Republicans to Democrats in response to unionization, with no effect on total contribution amounts.
@@
Effects of losing a union election. Our DiD results measure the differential change in contributions from establishments where the union won versus establishments where the union lost the election. The observed relative shift in donations could not only be explained by the effects of unionization after winning the election but also by the effect of holding and losing an election. Interaction with the union organization in preparation for the union election as well as a potentially increased salience of worker issues and distributional conflicts may affect the political behavior of employees, in particular in the short term, even if the union election is lost. We test this by estimating the effects of losing an election compared to holding no election. To avoid selection into which establishments hold and lose elections, we exploit only variation in the timing of union elections and use establishments that hold and lose an election in the future as the control group. We implement this approach in a stacked DiD model similar to our baseline model (1). [^36] See Appendix C for details of the stacking implementation. We also run the DiD estimators of Borusyak et al. (forthcoming) and Callaway and Sant’Anna (2021), which yield similar results. Results are presented in Appendix Table A.9. We obtain small and insignificant estimates for our two main outcomes and for both workers and managers with a precision similar to our baseline results. This suggests that losing a union election can indeed be viewed as an untreated counterfactual and that our results are driven by the effect of unionization after winning a union election.
Overall, our estimates provide robust evidence that unionization changes the composition of employees’ campaign contributions in favor of Democratic (relative to Republican) candidates. Importantly, this effect is found for both workers and managers.
@@ 6.1 Composition versus Individual-Level Effects
We now examine whether the establishment-level effects of unionization on employees’ campaign contributions indeed reflect changes in contributions of individual employees. Alternatively, the total contribution amount of an establishment may be affected by a change in the number of employees – which, unfortunately, we do not have information on – and the partisan leaning of employees’ contributions may be driven by compositional changes regarding what type of employees separate from and are newly hired into unionized establishments. Frandsen (2021) finds that unionization leads older and higher-paid workers to leave and younger workers to join union jobs. Separations and hirings may also be selective in terms of political ideology. For example, conservative union-avoiding managers may want to leave unionized workplaces and be replaced by more liberal ones. To disentangle compositional and individual-level changes, we exploit donor identifiers in the DIME, which allows us to track donors’ contributions over time.

### v5

@@ 5.2 Addressing Identification Challenges
Vote share tests. We continue presenting results for our RDD-motivated tests to probe the validity of the underlying parallel trends assumption of the DiD model. [^33] One particular concern for the parallel trends assumption would arise if union elections were endogenously timed around federal election dates. Appendix Figure A.3 investigates whether union elections follow political cycles. Across years with and without federal elections, there are no strong differences in the number of union elections held and the probability of winning a union election, in particular not around the week of federal elections. Thus, we do not see evidence that employers or unions successfully manipulate union election dates to change union support around federal election cycles. Figure 3 focuses on the partisan composition of contributions, while effects on the total amount of contributions are presented in Appendix Figure A.4. Results are always reported separately for workers and managers. We first analyze the heterogeneous effects of unionization across the vote share distribution. Panel (a) of Figure 3 displays the $\delta_{g}$ coefficients from model (3) for the interaction between the post-election dummy and different vote share categories. The results show that there are no significantly different trends among losing elections with a vote share of 0-20% or 35-50% relative to those with 20-35%, for contributions from both workers and managers. The partisan composition of contributions thus appears to evolve similarly across losing establishments with different vote shares, suggesting that unobserved shocks correlated with voting results are unlikely to drive our results. [^34] In Appendix Figure A.5, we also investigate whether pre-trends in the contribution composition are similar across the vote share distribution. For this, we estimate the following modified version of model (3): $y_{jk}=\alpha_{j}+\beta_{kg_{j}}+\sum_{g}\delta^{PRE}_{g}\times\Big(\mathbbm{1}[k<-1]\times\mathbbm{1}[V_{j}\in\nu^{g}]\Big)+\sum_{g}\delta^{POST}_{g}\times\Big(\mathbbm{1}[k\geq 0]\times\mathbbm{1}[V_{j}\in\nu^{g}]\Big)+\epsilon_{jk}$ (4) The results show that none of the $\delta^{PRE}_{g}$ coefficients are significantly different from zero, indicating that, also before the union election, contribution patterns evolved similarly across establishments with different voting results. Moreover, the results indicate whether treatment effects are heterogeneous across vote shares among won union elections. For the composition of contributions from managers, the estimate is significant across all vote share categories above 50%. Thus, the political response of managers does not appear to depend on whether workers won the union election with large or small margins of victory. For workers, the effect on partisan support is significant only for vote shares between 50 and 80% and appears smaller for elections won by a large margin.
@@
We further examine the robustness of our DiD estimates when focusing on close union elections. Establishments with narrowly decided votes are likely to be more similar in characteristics and to be exposed to similar shocks, making the parallel trends assumption more credible. Panel (b) of Figure 3 presents coefficients from model (1), estimated on subsamples restricted to increasingly narrow union vote share bandwidths (in 5% steps) around the 50% cutoff. Our baseline results from Figure 2 include only elections with a pro-union vote share between 20 and 80%, i.e., a bandwidth of 30%. Figure 3 shows that treatment effects are very similar when instead using all elections. Importantly, the results are also very stable when focusing on closer elections. Even when restricting the sample to establishments that won with a maximum vote margin of 5%, we see a positive and significant effect on the composition of campaign contributions for managers. Similarly, for workers a maximum vote margin of 10% already yields a positive and significant effect.
Alternative source of variation. To further check the causal interpretation of our results, we also complement the DiD strategy with a novel IV approach. For this, we exploit variation in unionization resulting from the timing of exogenous shocks to the salience of safety at work that are triggered by unexpected fatal workplace accidents shortly before the union election. After the NLRB accepts a petition to hold a union election, it sets the timeline of the unionization process and fixes an election date. Any random unexpected shocks between petition and election that shift union support are potential candidates for an instrument. We focus on sector-level fatal work accidents in the 30 days before a union election. Unions often campaign on safety issues and are found to improve safety conditions in the workplace (e.g., AFLCIO2022; Hagedorn2016; Li2022). An increase in the salience of serious work-safety issues thus plausibly strengthens union support in the union election.
The exclusion restriction of the instrument relies on the notion that a shock in fatal work accidents in the same sector affects political behavior in subsequent years only through its impact on the likelihood that an establishment will unionize. Two points are worth highlighting in that regard. First, while all individuals in our sample are potentially exposed to the information on fatal work accidents, only some vote on unionization in the following 30 days. We control for accidents in the same sector and year, such that we only exploit variation in the timing of the information shocks relative to the union election. Second, we focus on the medium-term impact of spikes in fatal accidents, excluding observations in the cycle of the union election. The result that common shocks in fatal work accidents influence political behavior years afterward in some but not other establishments would be difficult to explain other than through the path dependency triggered by the increase in the likelihood of unionization shortly after the accidents.
Appendix C describes the implementation of the IV approach and presents the results. While the IV estimates confirm our main finding – that unionization leads to a leftward shift in campaign contributions – they are less precise than the DiD estimates. We therefore consider the IV approach as a supplementary strategy that lends additional support to our main results.
@@ 5.3 Robustness
We now discuss further robustness checks for our main DiD estimates. Results are presented in Appendix Tables A.5 to A.9 and D.1.
Extensive versus intensive margin weights. The percentage effect interpretation of the IHS results implicitly assumes a weight of extensive and intensive margin changes in contribution amounts (Chen2024). To understand the role of both margins, we perform two analyses. First, in Appendix Table A.5, we directly study extensive margin effects using as outcome variable a dummy for any positive donation by an employee at the establishment. The DiD estimates show no significant effect of unionization on the likelihood that at least one employee in the establishment donates a positive amount to any candidate. However, for the partisan composition, we find that unionization results in an increased likelihood of employees donating to Democratic relative to Republican candidates of 2.8 percentage points. This effect is observed for both workers’ and managers’ donations and suggests that extensive margin changes may explain at least some of the partisan shift in campaign contributions.
Second, we analyze the robustness of our main results when placing explicit weights on extensive and intensive margin changes. For that, we follow the suggestion in Chen2024 and use the following outcome transformation:
$m(\mathdollar)=\begin{cases}\ln(\mathdollar)&\text{for }\mathdollar>0\\ -x&\text{for }\mathdollar=0.\end{cases}$ (5)
This transformation assigns equal weight to the extensive margin effect of moving from 0 to 1 and the intensive margin effect of increasing donations by 100xx log points (approximately 100xx% percent) for establishments with positive donations. Appendix Table A.6 reports results for alternative weights $x\in\{0,0.1,0.7,1,3\}$. Setting $x=0.7$, we find the same DiD estimates as for the IHS-transformed outcomes, which indicates that our main results implicitly weight an extensive margin change from 0 to 1 as equivalent to an intensive margin change in donations by 70 log points. [^35] This equivalence is expected, given that $\lim_{y\to\infty}\left(\text{arsinh}(y)-\ln(y)\right)=\lim_{y\to\infty}\left(\ln\left(y+\sqrt{y^{2}+1}\right)-ln(y)\right)=\ln(2)\approx 0.7.$ Importantly, the results remain similar, growing only slightly in magnitude, for alternative weights of 10, 100, or 300 log points. For $x=0$, which shuts off extensive margin changes by treating zero amounts as equal to one USD, we also find very similar results. Across all weights, we estimate increases in the Democratic-Republican gap between 11 and 16 (19 and 26) percentage points among workers (managers).
@@
Outcome transformations.Roth2021 point out that different transformations of the outcome variable may imply different parallel trends assumptions in a DiD design. We therefore test the sensitivity of our results to alternative outcome transformations. In addition to transformations using IHS and equation (5), we consider the log function after adding one to the amounts, take the quartic root of amounts [^36] As shown by Thakral2023, power transformations – unlike quasi-logarithmic transformations (such as the IHS or log(y+1) function) – maintain scale invariance of the implied semi-elasticities., as well as leave amounts untransformed (in 2010 USD). Results, shown in Panels B, C, and D of Appendix Table A.7, yield the same conclusions as the results for the IHS-transformed outcomes.
@@
Staggered DiD estimators. The recent econometrics literature has proposed different methods to circumvent issues of treatment effect heterogeneity in staggered DiD designs. All the proposed estimation strategies have in common that they restrict the set of effective comparison units by ruling out the use of early-treated units in the estimation of treatment effects for currently-treated units. They differ, however, in terms of how exactly comparison units are identified and used in the estimation, as well as in terms of how cohort- or individual-specific treatment effect estimates are aggregated. [^37] In our stacking approach of model (1), we effectively only compare won elections to lost elections that were held in the same period, i.e., we only use never-treated units in the comparison group. The strategies by Borusyak2021 and Callaway2021, in contrast, also allow including not-yet-treated units in the comparison group. Both approaches differ in that Borusyak2021 use the average pre-treatment outcome over all pre-treatment periods, whereas Callaway2021 only use the outcome one period before treatment start. In terms of aggregation, Gardner2021 shows that the stacking approach identifies a convexly weighted average of cohort-specific treatment effects where the weights are given by the number of treated units and the variance of treatment within each cohort. In comparison, Borusyak2021 and Callaway2021 first estimate unit- or cohort-specific effects and then aggregate through a simple average across treated units. Callaway2021 also allow other weights, but we use the default option where cohort-specific estimates are weighted by the number of treated units in each cohort. In Panels E and F of Appendix Table A.7, we present results from the imputation approach of Borusyak2021 and the estimator developed by Callaway2021. The estimates are very similar to our stacked DiD results.
@@
Sample restrictions. For our baseline results, we have restricted the sample to establishments with at least one matched contribution over the observation period. If employees have not contributed before the union election and start contributing in response to unionization, concerns of sample selection bias may arise. In Appendix Table A.2, we show that there are no effects of unionization on the likelihood of observing any contribution after the union election. Nevertheless, Panel G of Appendix Table A.7 also shows results when restricting the analysis to establishments that have at least one matched contribution in any cycle before the union election. This reduces our sample size but leaves results largely unchanged. Another concern may arise from selected establishment survival (Frandsen2020; Wang2021). In particular, we may wrongly assign zero contribution amounts to establishments that do not yet exist or that have already been closed. In order to restrict the analysis to a window of known survival, we exclude for each establishment the cycles before the first matched contribution and the cycles after the last matched contribution. As shown in Panel H, this restriction increases the size of our estimates suggesting that our baseline results can be interpreted as lower bounds.
@@
Manager-worker classifications. In Appendix Table A.8, we check whether our results are sensitive to the exact definition of managers and supervisors versus non-managerial workers. To see whether the political response is different for lower- and upper-tier managers, we use more stringent definitions of managers/supervisors. First, we vary the cutoff for the importance of supervisor tasks (Panels B and C). Second, we only consider “Management Occupations” (SOC group 11) and treat all other occupations (including those with high importance of supervisor tasks) as workers (Panel D). The results do not change much with these alternative classifications. Even for more upper-tier managers we find an increase in the support for Democrats over Republicans. As the contribution data contains a non-negligible share of occupations that we were unable to classify, we also use donors’ exact residence location as an alternative predictor of occupational status that is available for the complete data. Specifically, we identify managers and workers by whether they live in a census tract with a median household income higher or lower than the 80th or 90th percentile of the state-specific distribution of census-tract incomes (Panels E and F). We consistently find positive effects of unionization on Democratic vs. Republican support for donors who live in high- and low-income neighborhoods.
@@
Federal versus state candidates. Next, we examine whether our effects are limited to contributions to candidates for either federal or state offices. Focusing on federal races offers the advantage of analyzing a set of comparable candidates who are potentially relevant to citizens nationwide. At the same time, U.S. legislation on labor issues, which unions may focus on when endorsing candidates, is enacted not only at the federal but also at the state level (e.g., state-specific minimum wages, right-to-work laws). Appendix Table A.9 shows that our estimates are driven by contributions to federal and state candidates. The effect sizes are somewhat larger for contributions to candidates running for federal office, but at both levels, we observe a significant shift in donations from Republicans to Democrats in response to unionization.
@@
Effects of losing a union election. Our DiD results measure the differential change in contributions between establishments with won and lost union elections. The observed relative shift in donations may reflect not only unionization effects after winning the election but also effects from holding and losing an election. Interactions with union organizers and heightened awareness of worker issues and distributional conflicts could affect employees’ political behavior, in particular in the short term, even if the union election is lost. To test this, we estimate the effects of losing an election compared to holding no election, leveraging variation in the timing of elections. Specifically, we implement a stacked DiD model, using as the control group establishments that hold and lose an election in the future (see Appendix D for details of the stacking implementation). Results, presented in Appendix Table D.1, show small and insignificant estimates for our two main outcomes and for both workers and managers, with a precision similar to our baseline results. This suggests that losing an election serves as a valid counterfactual, confirming that our results capture the effects of unionization after winning a union election.
@@ 6.1 Composition versus Individual-Level Effects
We now examine whether the establishment-level effects of unionization on employees’ campaign contributions indeed reflect changes in contributions of individual employees. Alternatively, the total contribution amount of an establishment may be affected by a change in the number of employees and the partisan leaning of employees’ contributions may be driven by compositional changes regarding what type of employees separate from and are newly hired into unionized establishments. Frandsen2020 shows that unionization leads older and higher-paid workers to leave and younger workers to join union jobs. Separations and hirings may also be selective in terms of political ideology. For example, conservative union-avoiding managers may want to leave unionized workplaces and be replaced by more liberal ones. To disentangle compositional and individual-level changes, we exploit donor identifiers in the DIME, which allows us to track donors’ contributions over time.